Tom, dick, and harry are camping in their tents. If the distance between tom and dick is 153 feet, the distance between tom and harry is 201 feet, and the distance between dick and harry is 175 feet, what are the angles between dick, harry and tom?
Answer by Alan3354(69443)

Tom, dick, and harry are camping in their tents. If the distance between tom and dick is 153 feet, the distance between tom and harry is 201 feet, and the distance between dick and harry is 175 feet, what are the angles between dick, harry and tom?
----------
Label the vertices T, D & H.
The sides opposite are t, d & h.
-----------
Use the Cosine Law to find a side.
t^2 = d^2 + h^2 - 2*d*h*cos(t)
-------------
I would use the Law of Sines to find a 2nd angle.
---
3rd angle = 180 - (sum of the other 2)